The sum of two numbers is 57. The smaller number is 13 less than the larger number. What are the two numbers?
Answer by zoomkaboom4(62)

2n-13= 57
+13 +13
2n= 70
/2 /2
n= 35
The numbers are 35 and 22.
Any more questions? Message me!
-Megan (Zoomkaboom4)